A football is approximately 1 foot long.
Football field is 120 yards, 360 feet long.
If I wanted to make a model of a football field that is 36cm long, how big would a football in this scale model have to be?

Answer by rfer(16322)

36/360=x/1
360x=36
x=36/360
x=0.1cm